# arrow size=1pt, every edge=line width=0.5pt

### v1

@@ 1 Introduction
The nucleon axial-vector coupling constant $g_{A}$ is a fundamental parameter of low-energy chiral QCD dynamics as it parameterizes the leading order pion-nucleon interaction. It further plays an eminent role in neutron $\beta$-decay, see Ref. [2] for a recent work, in nuclear reactions and in neutrino scattering off nuclei, as witnessed by the T2K, NOvA, MINERvA, MicroBooNE, and SBN experiments. For reviews, see Refs. [3, 4]. Furthermore, the axial-vector coupling is also considered a “gold-plated” observable for the ab initio lattice QCD approach. The first high-precision lattice QCD calculation was reported in Ref. [5] and a summary of older and more recent results is collected in Ref. [6]. Most of these simulations are performed for unphysical pion masses, while in Refs. [7, 8, 9, 10] the physical pion mass is also considered. In any case, the issue of precise and controlled chiral extrapolations is still of relevance. Of course, it is also of general interest to study the higher-order corrections to this observable, as they encode information about the convergence of two-flavor baryon chiral perturbation theory. Therefore, in this work, we consider the calculation of $g_{A}$ to two loops using two different covariant schemes, namely the extended-on-mass-shell (EOMS) approach [11] and the method of infrared regularization (IR) [12]. Detailed discussions of these schemes and comparisons with the often used heavy baryon approach can be found in Refs. [13, 14]. This work differs from the earlier paper [15], where renormalization group methods were applied to the chiral pion-nucleon Lagrangian in the heavy baryon approach and the two-loop representation was confronted with then existing lattice calculations at unphysical pion masses. We will come back to that work later.
@@ 2 Chiral expansion of $g_{A}$
In baryon chiral perturbation theory (BCHPT), one encounters odd and even powers of the small expansion parameter $q$, which in the two flavor case is given by pion masses and momenta as well as nucleon three-momenta. The nucleon mass $m_{N}$ is of the same size as the hard scale related to chiral symmetry breaking, often estimated as $\Lambda_{\chi}=4\pi F_{\pi}\simeq 1.2\,$GeV, with $F_{\pi}\simeq 92\,$MeV the pion decay constant. Tree diagrams are of order $q$ and $q^{2}$, one-loop diagrams of order $q^{3}$ and $q^{4}$, and the two-loop corrections are of order $q^{5}$ and $q^{6}$, respectively. Consequently, the chiral expansion of the axial-vector coupling takes the form
@@
$\quad+\left(\frac{\alpha_{4}}{(4\pi F)^{4}}\ln^{2}\frac{M}{\mu}+\frac{\gamma_{4}}{(4\pi F)^{4}}\ln\frac{M}{\mu}+\beta_{4}\right)\,M^{4}+\alpha_{5}\,M^{5}\biggr{\}}+{\cal O}(M^{6})~{},$
$=$ $g_{0}\,\,\biggl{\{}1+\underbrace{\Delta^{(2)}+\Delta^{(3)}}_{\rm one-loop}+\underbrace{\Delta^{(4)}+\Delta^{(5)}}_{\rm two-loop}\biggr{\}}+{\cal O}(M^{6})~{},$
@@
$g_{A}=g_{0}\,\left[1+{\cal O}(M^{2})\right]\,\,.$ (2)
Further, $\mu$ is the scale of dimensional regularization and $\Delta^{(n)}$ represents the correction at order $n$ in the chiral counting. We have expressed $g_{A}$ in terms of the pion decay constant in the chiral limit, $F$, and the leading order term of the quark mass expansion of the pion mass squared, $M^{2}$. One has:
@@
The difference between $F$ and $F_{\pi}$ ($M$ and $M_{\pi}$) is of higher order when working at ${\cal O}(q^{3})$, however, at the order we are working it has to be taken into account. The third order one-loop coefficients $\alpha_{2}$ and $\beta_{2}$ in Eq. (1) were first given in Ref. [16] and the one-loop fourth order calculation was completed in Ref. [17] with (see also Ref. [18]),
@@
$g_{0}=g_{0}+\delta g_{0}|_{\rm reg}\,\quad d_{16}=d_{16}+\delta d_{16}|_{\rm reg}~{},$ (5)
@@
We now turn to the two-loop corrections. Note that contrary to the work [15], in which the coefficient $\alpha_{4}$ was calculated in heavy baryon chiral perturbation theory using renormalization group methods, we use here the operator basis and low-energy constants (LECs) given in Ref. [20]. At two-loop order there are four types of contributions: the irreducible diagrams, the one coming from the wave function and coupling constant renormalization, the one loop graphs with vertices from ${\cal L}_{\pi N}^{(3)}$, and counterterms. One thus has schematically
$g_{A}^{\rm 2-loop}\equiv g_{0}\,\left(\Delta^{(4)}+\Delta^{(5)}\right)=g_{A}^{\rm ren}+g_{A}^{d_{i}}+g_{A}^{\rm irr}+g_{A}^{\rm ct}~{}.$ (6)
In this paper we will concentrate on the $\Delta^{(4)}$ contributions, meaning that we give here the $M^{4}$ contribution to $g_{A}$, the full two-loop order calculation is in progress [21]. The latter involves the calculation of several two-loop diagrams with vertices from ${\cal L}_{\pi N}^{(2)}$.
@@ 3 Outline of the calculation
Figure 1: Two-loop diagrams contributing to the nucleon self-energy. The same topologies contribute to $g_{A}$. For the latter one has to hook the axial current wherever possible on the nucleon propagator and on the vertices with pions and nucleons. The dashed lines represent the pion and the solid ones the nucleon. The dots are vertices from ${\cal L}_{\pi N}^{(1)}$ at the order we are working.
@@
with $\Sigma(p^{2})$ the nucleon self-energy, which has a similar expansion as discussed above for $g_{A}$, namely
@@
Figure 2: Two-loop diagrams contributing to $g_{A}$. The wavy line represents the axial current, the dashed ones the pion and the solid ones the nucleon.
@@
with the $\delta_{i}$ given in Ref. [20] for the IR case and in Ref. [19] for the full EOMS case. Note that each $\delta_{i}$ can contribute to various powers of $g_{0}$ as one has
$\delta_{i}=\delta_{i}^{(0)}+\delta_{i}^{(1)}g_{0}+\delta_{i}^{(2)}g_{0}^{2}+\delta_{i}^{(3)}g_{0}^{3}+\delta_{i}^{(4)}g_{0}^{4}\,,$ (19)
where the $\delta_{i}$ with $i=1,2$ are contributing only to even powers of $g_{0}$, $\delta_{14}$ contributes only to $g_{0}^{4}$ and the others only to odd powers of $g_{0}$. This has to be kept in mind when checking the scale dependence. There are also contributions from the mesonic LECs $l_{3}$ and $l_{4}$, which enter the quark mass expansion of the pion mass and the pion decay constant, respectively. Similarly to the $d_{i}$ one has
@@
where $d_{i}|_{IR}$ contains the $\delta_{i}$ from Ref. [20] and $d_{i}|_{R}$ contains the difference between the EOMS and IR results for the $\delta_{i}$. The first sum in the last equality in Eq. (22) added to the IR two loop contributions constitutes the IR result. It satisfies the Ward identities. The same is true for the second sum together with the mixed terms from the two-loop graphs, which we will call the mixed result. Finally one has the contribution from the third sum and the regular part of the two loop graphs which can be absorbed into $g_{A}^{\rm ct}$. The EOMS result is the sum of these three parts.
@@
which is appropriate for the modified $\overline{\rm MS}$ subtraction scheme used here as it is customary in CHPT (for similar results in two-loop calculation in the meson sector see, e.g., Ref. [34]). Let us consider the derivative of $C$ with respect to $\mu$. It turns out that the contribution to $g_{0}^{i}$ of this derivative is not given by the derivative of $C^{(i)}$ with respect to $\mu$ as the latter can contribute to various power of $g_{0}$. This is specific to the baryon sector as already encountered in the case of the $d_{i}$, see discussion after Eq. (19). The $C^{(i)}$ therefore can not be individually scale independent. Thus one has
@@
To get these relations provides another check of the calculation together with the scale independence of the result. One obtains
@@ 4 Results and discussion
$\gamma_{4}^{(1)}$ $=$ $-32\pi^{2}F^{2}\,\left(d_{14}^{r}-2(d_{1}^{r}+d_{2}^{r})\right)-64\pi^{2}\left(l_{3}^{r}-l_{4}^{r}\right)-\frac{389}{36}~{},$
@@
$\gamma_{4}^{(3)}$ $=$ $-128\pi^{2}\,\left(l_{3}^{r}-l_{4}^{r}\right)+\frac{1}{9}\left(13-16\pi^{2}\right)~{},$
@@
And finally one has for $\beta_{4}$
@@
$(4\pi F)^{4}\beta_{4}^{(1)}$ $=$ $-8\pi^{2}\,\left(d_{14}^{r}+2(d_{1}^{r}+d_{2}^{r})\right)-32\pi^{2}l_{3}^{r}+\frac{3575}{864}-\frac{\pi^{2}}{3}+\frac{1}{2}\psi^{(1)}\biggl{(}\frac{2}{3}\biggr{)}+C_{0}^{(1)}~{},$
@@
$(4\pi F)^{4}\beta_{4}^{(3)}$ $=$ $32\pi^{2}\,\left(-3l_{3}^{r}+l_{4}^{r}\right)-\frac{\pi^{2}}{27}\left(61+48\log 3\right)-\frac{335}{432}-\frac{1}{6}\psi^{(1)}\biggl{(}\frac{2}{3}\biggr{)}+C_{0}^{(3)}~{},$
@@
where $\psi^{(1)}(2/3)=3.06388$ is the first derivative of the digamma function at $2/3$. Note that the $\beta_{4}^{(i)}$ also have contributions from the $d_{i}^{\epsilon}$ and $l_{i}^{\epsilon}$ terms in Eqs. (18) and (20), respectively. These can be absorbed in the $C_{0}^{(i)}$.
@@
At the next order there will be $1/m$ corrections to these results. In the EOMS scheme the mixed terms do not contribute to $\alpha_{4}^{(i)}$. This has to be the case as the leading non-analytic terms have to be the same in all renormalization schemes. This is in principle not the case for $\gamma_{4}^{(i)}$ as this coefficient is renormalization scheme-dependent due to the different treatment of one-loop diagrams in different renormalization schemes. It turns out that in our case the mixed terms do not contribute to $\gamma_{4}^{(i)}$. Thus the only difference between EOMS and IR at that order are terms contributing to $\beta_{4}$. These will be hidden in our $C_{0}$’s.
@@
Let us now discuss the convergence of the series and the dependence of $g_{A}$ on the pion mass which is relevant for the lattice. We will not perform fits to lattice data here, but rather use typical values for the pertinent LECs from various analyses of elastic and inelastic pion-nucleon scattering in the EOMS scheme to get an idea about the size of the leading two-loop corrections. A more detailed discussion including also an uncertainty analysis will be given in Ref. [21]. To be concrete, we use $F_{\pi}=0.927\,$GeV and $M_{\pi}=0.139\,$GeV, and the nucleon mass in the chiral limit is taken as $m=0.87\,$GeV [36]. Further, we set the renormalization scale to $\mu=m$ (which leads to $\log(m/\mu)$=0, a quantity which appears in the mixed terms and in the regular ones in principle) and take the values of the LECs at that scale. The LECs in the mesonic sector are rather well known, one has $l_{3}(m)=1.4\cdot 10^{-3}$ and $l_{4}(m)=3.7\cdot 10^{-3}$. From the values of $l_{3}$ and $l_{4}$ one can determine $F$ and $M$. In the baryon sector the LECs are less known. From an analysis of elastic and inelastic pion-nucleon scattering [37] we take set 1 (which is based on the standard power counting $m_{N}\sim\Lambda_{\chi}$) $c_{3}=-6.63\,$GeV-1, $c_{4}=4.01\,$GeV-1, $\bar{d}_{1}+\bar{d}_{2}=4.37\,$GeV-2, $\bar{d}_{10}=-0.8\,$GeV-2, $\bar{d}_{11}=-15.6\,$GeV-2, $\bar{d}_{12}=5.9\,$GeV-2, $\bar{d}_{13}=13.6\,$GeV-2, $\bar{d}_{14}=-7.43\,$GeV-2, and $\bar{d}_{16}=0.4\,$GeV-2, and as set 2 (which is based on the power counting used in nucleon-nucleon scattering $m_{N}\sim\Lambda_{\chi}^{2}/M_{\pi}$), $c_{3}=-7.26\,$GeV-1, $c_{4}=4.74$ GeV-1, $\bar{d}_{1}+\bar{d}_{2}=3.39,$GeV-2, $\bar{d}_{10}=10.9\,$GeV-2, $\bar{d}_{11}=-30.9\,$GeV-2, $\bar{d}_{12}=-10.9\,$GeV-2, $\bar{d}_{13}=27.7,$GeV-2, $\bar{d}_{14}=-7.36\,$GeV-2, and $\bar{d}_{16}=-3.0\,$GeV-2. Note that we only have information on $\bar{d}_{14}-\bar{d}_{15}$ and we have assumed here $\bar{d}_{15}=0$. Further, the $\bar{d}_{i}$ are the $d_{i}^{r}(\mu)$ defined at $\mu=M_{\pi}$, see, e.g., Ref. [38]. These central values have been obtained in heavy baryon chiral perturbation theory from a combined fit to the reactions $\pi N\to\pi N$ and $\pi N\to\pi\pi N$. As we have done an expansion in $M/m$ of the IR and EOMS expressions our results should in fact correspond to the one in the heavy baryon approach. Finally, $\bar{d}_{18}=-0.8\,$GeV-2 can be related to the Goldberger-Treiman discrepancy [38, 39]. For the LECs $C_{0}^{(i)}$ we will assume that they are of natural size, typically of the order of $1/\Lambda^{4}_{\chi}$ with $\Lambda_{\chi}=0.6\,$GeV an estimate of the breakdown scale of the chiral expansion (which is a more conservative estimate than given above and was used in Refs. [19, 37]). In Fig. 3 (left panel) we show the chiral expansion of $g_{A}$ at second, third and fourth order for set 1 of the LECs and similarly for set 2 in the right panel of Fig. 3. The upper fourth order curve corresponds to $C_{0}^{(i)}(m)=15\,$GeV-4, whereas the lower fourth order curve represents the case $C_{0}^{(i)}(m)=-15\,$GeV-4. While the third order correction is large, as the coefficient $\alpha_{3}$ in Eq. (1) is proportional to the large factor $2c_{4}-c_{3}$, we see that the fourth order corrections are rather small for pion masses below 300 MeV, even though some of the dimension-three LECs are rather large. Also, we find that $g_{0}=1.0$ for set 1 and $g_{0}=1.3$ for set 2, in order.
@@
Figure 3: $g_{A}$ as a function of $M$ for set 1 (left panel) and set 2 (right panel) of the LECs given in the text. The blue dash-dotted line, the black dashed line and the red solid lines represent the results at order $q^{2}$, $q^{3}$ and $q^{4}$, respectively. The upper $q^{4}$ curve corresponds to $C_{0}^{(i)}(m)=15\,$GeV-4, whereas the lower $q^{4}$ curve represents the case $C_{0}^{(i)}(m)=-15\,$GeV-4. The pink circle denotes the physical value of $g_{A}$.
@@
${\rm set}~{}1$ ::: $\Delta^{(2)}=1.5\%~{},\,~{}~{}~{}\quad\Delta^{(3)}=28.8\%~{},\quad\Delta^{(4)}=3.9\%~{},$
${\rm set}~{}2$ ::: $\Delta^{(2)}=-26.7\%~{},\quad\Delta^{(3)}=44.5\%~{},\quad\Delta^{(4)}=-15.3\%~{},$ (34)

### v2

@@ 1 Introduction
The nucleon axial-vector coupling constant $g_{A}$ plays an eminent role in neutron $\beta$-decay, see Ref. [2] for a recent work, in nuclear reactions and in neutrino scattering off nuclei, as witnessed by the T2K, NOvA, MINERvA, MicroBooNE, and SBN experiments. For reviews, see Refs. [3, 4]. Furthermore, $g_{A}$ is also a fundamental parameter of low-energy chiral QCD dynamics which parameterizes the leading order pion-nucleon interaction as at tree level the Goldberger-Treiman relation (GTR) allows to express the pion-nucleon coupling in terms of the axial-vector coupling. Corrections to the GTR are known to be small so that this relation is an approximate one when considering loops. Furthermore, the axial-vector coupling is also considered a “gold-plated” observable for the ab initio lattice QCD approach. The first high-precision lattice QCD calculation was reported in Ref. [5] and a summary of older and more recent results is collected in Ref. [6]. Most of these simulations are performed for unphysical pion masses, while in Refs. [7, 8, 9, 10] the physical pion mass is also considered. In any case, the issue of precise and controlled chiral extrapolations is still of relevance. Of course, it is also of general interest to study the higher-order corrections in the quark mass expansion of this observable, as they encode information about the convergence of two-flavor baryon chiral perturbation theory. Therefore, in this work, we consider the calculation of $g_{A}$ to two loops using two different covariant schemes, namely the extended-on-mass-shell (EOMS) approach [11] and the method of infrared regularization (IR) [12]. Detailed discussions of these schemes and comparisons with the often used heavy baryon approach can be found in Refs. [13, 14]. This work differs from the earlier paper [15], where renormalization group methods were applied to the chiral pion-nucleon Lagrangian in the heavy baryon approach and the two-loop representation was confronted with then existing lattice calculations at unphysical pion masses. We will come back to that work later.
@@ 2 Chiral expansion of $g_{A}$
In baryon chiral perturbation theory (BCHPT), one encounters odd and even powers of the small expansion parameter $q$, which in the two flavor case is given by pion masses and momenta as well as nucleon three-momenta. The nucleon mass $m_{N}$ is of the same size as the hard scale related to chiral symmetry breaking, often estimated as $\Lambda_{\chi}=4\pi F_{\pi}\simeq 1.2\,$GeV, with $F_{\pi}\simeq 92\,$MeV the pion decay constant. Tree diagrams start contributing at order $q$, one-loop diagrams at order $q^{3}$, two-loop ones at order $q^{5}$, three-loop diagrams at order $q^{7}$, and so on. In this paper we will concentrate on the leading two-loop diagrams of chiral order $q^{5}$, the order $q^{6}$ two-loop calculation is in progress [21]. The latter involves the calculation of several two-loop diagrams with vertices from ${\cal L}_{\pi N}^{(2)}$. Consequently, the chiral expansion of the axial-vector coupling as evaluated in this work takes the form
@@
$\quad+\left(\frac{\alpha_{4}}{(4\pi F)^{4}}\ln^{2}\frac{M}{\mu}+\frac{\gamma_{4}}{(4\pi F)^{4}}\ln\frac{M}{\mu}+\beta_{4}\right)\,M^{4}\biggr{\}}+{\cal O}(M^{5})~{},$
$=$ $g_{0}\,\,\biggl{\{}1+\underbrace{\Delta^{(2)}+\Delta^{(3)}}_{\rm one-loop}+\underbrace{\Delta^{(4)}}_{\rm leading~{}two-loop}\biggr{\}}+{\cal O}(M^{5})~{},$
@@
$g_{A}=g_{0}\,\left[1+{\cal O}(M^{2})\right]\,,$ (2)
and $\Delta^{(4)}$ contains the two loop contribution of chiral order $q^{5}$, as well as contributions from one loop diagrams with vertices from ${\cal L}_{\pi N}^{(1,2,3)}$. Note that terms with denominations one-loop/leading two-loop in the last line of Eq. (1) include the contributions from tree diagrams with vertices from ${\cal L}_{\pi N}^{(3/5)}$. Further, $\mu$ is the scale of dimensional regularization and $\Delta^{(n)}$ represents the correction at order $n$ in the chiral counting. We have expressed $g_{A}$ in terms of the pion decay constant in the chiral limit, $F$, and the leading order term of the quark mass expansion of the pion mass squared, $M^{2}$. One has:
@@
The difference between $F$ and $F_{\pi}$ ($M$ and $M_{\pi}$) is of higher order when working at ${\cal O}(q^{3})$, however, at the order we are working it has to be taken into account. The one-loop coefficients $\alpha_{2}$ and $\beta_{2}$ in Eq. (1) were first given in Ref. [16] and the one-loop fourth order calculation was completed in Ref. [17] with (see also Ref. [18]),
@@
$g_{0}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\to}\,g_{0}+\delta g_{0}|_{\rm reg}\,\quad d_{16}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\to}\,d_{16}+\delta d_{16}|_{\rm reg}~{},$ (5)
@@
Let us now discuss the $M^{4}$ contribution to $g_{A}$. Note that contrary to the work [15], in which $\alpha_{4}$ was calculated in heavy baryon chiral perturbation theory using renormalization group methods, we use here the operator basis and low-energy constants (LECs) given in Ref. [20]. There are four types of contributions to $\Delta^{(4)}$: the irreducible diagrams, the one coming from the wave function and coupling constant renormalization, the one loop graphs with vertices from ${\cal L}_{\pi N}^{(3)}$, and counterterms. One thus has schematically
$g_{0}\,\Delta^{(4)}=g_{A}^{\rm ren}+g_{A}^{d_{i}}+g_{A}^{\rm irr}+g_{A}^{\rm ct}~{}.$ (6)
@@ 3 Outline of the calculation
Figure 1: Two-loop diagrams contributing to the nucleon self-energy. The same topologies contribute also to $g_{A}$. For the latter one has to hook the axial current wherever possible on the nucleon propagator and on the vertices with pions and nucleons. The dashed lines represent the pion and the solid ones the nucleon. The dots are vertices from ${\cal L}_{\pi N}^{(1)}$ at the order we are working. Note that there are further two-loop diagrams shown in Fig. 2 where the axial-current couples to three pions.
@@
[…]
@@
Figure 2: Two-loop diagrams contributing to $g_{A}$. The wavy line represents the axial current, the dashed ones the pion and the solid lines the nucleon.
@@
with the $\delta_{i}$ given in Ref. [20] for the IR case and in Ref. [19] for the full EOMS case. Note that each $\delta_{i}$ has an expansion in powers of $g_{0}$, which up to the order we are working has the form
$\delta_{i}=\delta_{i}^{(0)}+\delta_{i}^{(1)}g_{0}+\delta_{i}^{(2)}g_{0}^{2}+\delta_{i}^{(3)}g_{0}^{3}+\delta_{i}^{(4)}g_{0}^{4}+{\cal O}(g_{0}^{5})\,,$ (19)
where the $\delta_{i}$ with $i=(1,2)$/$(10-13,16)$ have non-vanishing contributions of only even/odd powers of $g_{0}$ respectively, while for $i=14$ only $\delta_{14}^{(4)}$ is different from zero. This has to be kept in mind when checking the scale dependence. There are also contributions from the mesonic LECs $l_{3}$ and $l_{4}$, which enter the quark mass expansion of the pion mass and the pion decay constant, respectively. Similarly to the $d_{i}$ one has
@@
where $d_{i}|_{IR}$ contains the $\delta_{i}$ from Ref. [20] and $d_{i}|_{R}$ contains the difference between the EOMS and IR results for the $\delta_{i}$. The first sum in the last equality in Eq. (22) added to the IR two-loop contributions constitutes the IR result. It satisfies the Ward identities. The same is true for the second sum together with the mixed terms from the two-loop graphs, which we will call the mixed result. Finally one has the contribution from the third sum and the regular part of the two loop graphs which can be absorbed into $g_{A}^{\rm ct}$. The EOMS result is the sum of these three parts.
@@
which is appropriate for the modified $\overline{\rm MS}$ subtraction scheme used here as it is customary in CHPT (for similar results in two-loop calculation in the meson sector see, e.g., Ref. [34]). Let us consider the derivative of $C$ with respect to $\mu$. It turns out that the contribution to $g_{0}^{i}$ of this derivative is not given by the derivative of $C^{(i)}$ with respect to $\mu$ as the latter can contribute to various powers of $g_{0}$. This is specific to the baryon sector as already encountered in the case of the $d_{i}$, see discussion after Eq. (19). The $C^{(i)}$ therefore can not be individually scale-independent. Thus one has
@@
To get these relations provides another check of the calculation together with the scale-independence of the result. One obtains
@@ 4 Results and discussion
$\gamma_{4}^{(1)}$ $=$ $-32\pi^{2}F^{2}\,\left(d_{14}^{r}-2(d_{1}^{r}+d_{2}^{r})\right)-64\pi^{2}\left(l_{3}^{r}-l_{4}^{r}\right)-\frac{389}{36}$
$\quad-\frac{64\pi^{2}F^{2}}{m}\biggl{(}c_{2}+c_{3}-c_{4}-\frac{1}{2m}\biggr{)}~{},$
@@
$\gamma_{4}^{(3)}$ $=$ $-128\pi^{2}\,\left(l_{3}^{r}-l_{4}^{r}\right)+\frac{1}{9}\left(13-16\pi^{2}\right)+\frac{48\pi^{2}F^{2}}{m^{2}}~{},$
@@
And finally one has
$\beta_{4}=\beta_{4}^{(0)}+\beta_{4}^{(1)}g_{0}+\beta_{4}^{(2)}g_{0}^{2}+\beta_{4}^{(3)}g_{0}^{3}+\beta_{4}^{(5)}g_{0}^{5}~{},$ (33)
with
@@
$(4\pi F)^{4}\beta_{4}^{(1)}$ $=$ $-8\pi^{2}\,\left(d_{14}^{r}+2(d_{1}^{r}+d_{2}^{r})\right)-32\pi^{2}l_{3}^{r}+\frac{3575}{864}-\frac{\pi^{2}}{3}+\frac{1}{2}\psi^{(1)}\biggl{(}\frac{2}{3}\biggr{)}+C_{0}^{(1)}$
$\quad+\frac{16\pi^{2}F^{2}}{m}\biggl{(}c_{2}+4c_{4}+\frac{2}{m}\biggr{)}~{},$
@@
$(4\pi F)^{4}\beta_{4}^{(3)}$ $=$ $32\pi^{2}\,\left(-3l_{3}^{r}+l_{4}^{r}\right)-\frac{\pi^{2}}{27}\left(61+48\log 3\right)-\frac{335}{432}-\frac{1}{6}\psi^{(1)}\biggl{(}\frac{2}{3}\biggr{)}+C_{0}^{(3)}+\frac{32\pi^{2}F^{2}}{m^{2}}~{},$
@@
where $\psi^{(1)}(2/3)=3.06388$ is the first derivative of the digamma function at $2/3$. Note that the $\beta_{4}^{(i)}$ also have contributions from the $d_{i}^{\epsilon}$ and $l_{i}^{\epsilon}$ terms in Eqs. (18) and (20), respectively. These can be absorbed in the $C_{0}^{(i)}$. In the EOMS scheme the mixed terms do not contribute to $\alpha_{4}^{(i)}$. This has to be the case as the leading non-analytic terms have to be the same in all renormalization schemes. This is in principle not the case for $\gamma_{4}^{(i)}$ as this coefficient is renormalization scheme-dependent due to the different treatment of one-loop diagrams in different renormalization schemes. It turns out that in our case the mixed terms do not contribute to $\gamma_{4}^{(i)}$. Thus the only difference between EOMS and IR at that order are terms contributing to $\beta_{4}$. These will be hidden in our $C_{0}$s.
@@
Let us now discuss the convergence of the series and the dependence of $g_{A}$ on the pion mass which is relevant for the lattice. We will not perform fits to lattice data here, but rather use typical values for the pertinent LECs from various analyses of elastic and inelastic pion-nucleon scattering in the EOMS scheme to get an idea about the size of the leading two-loop corrections. A more detailed discussion including also an uncertainty analysis will be given in Ref. [21]. To be concrete, we use $F_{\pi}=0.927\,$GeV and $M_{\pi}=0.139\,$GeV, and the nucleon mass in the chiral limit is taken as $m=0.87\,$GeV [36]. Further, we set the renormalization scale to $\mu=m$ (which leads to $\log(m/\mu)$=0, a quantity which appears in the mixed terms and in the regular ones in principle) and take the values of the LECs at that scale. The LECs in the mesonic sector are rather well known, one has $l_{3}(m)=1.4\cdot 10^{-3}$ and $l_{4}(m)=3.7\cdot 10^{-3}$. From the values of $l_{3}$ and $l_{4}$ one can determine $F$ and $M$. In the baryon sector the LECs are less known. From an analysis of elastic and inelastic pion-nucleon scattering [37] we take set 1 (which is based on the standard power counting $m_{N}\sim\Lambda_{\chi}$) $c_{2}=3.51\,$GeV-1, $c_{3}=-6.63\,$GeV-1, $c_{4}=4.01\,$GeV-1, $\bar{d}_{1}+\bar{d}_{2}=4.37\,$GeV-2, $\bar{d}_{10}=-0.8\,$GeV-2, $\bar{d}_{11}=-15.6\,$GeV-2, $\bar{d}_{12}=5.9\,$GeV-2, $\bar{d}_{13}=13.6\,$GeV-2, $\bar{d}_{14}=-7.43\,$GeV-2, and $\bar{d}_{16}=0.4\,$GeV-2, and as set 2 (which is based on the power counting used in nucleon-nucleon scattering $m_{N}\sim\Lambda_{\chi}^{2}/M_{\pi}$), $c_{2}=4.89\,$GeV-1, $c_{3}=-7.26\,$GeV-1, $c_{4}=4.74$ GeV-1, $\bar{d}_{1}+\bar{d}_{2}=3.39,$GeV-2, $\bar{d}_{10}=10.9\,$GeV-2, $\bar{d}_{11}=-30.9\,$GeV-2, $\bar{d}_{12}=-10.9\,$GeV-2, $\bar{d}_{13}=27.7,$GeV-2, $\bar{d}_{14}=-7.36\,$GeV-2, and $\bar{d}_{16}=-3.0\,$GeV-2. Note that we only have information on $\bar{d}_{14}-\bar{d}_{15}$ and we have assumed here $\bar{d}_{15}=0$. Further, the $\bar{d}_{i}$ are the $d_{i}^{r}(\mu)$ defined at $\mu=M_{\pi}$, see, e.g., Ref. [38]. These central values have been obtained in heavy baryon chiral perturbation theory from a combined fit to the reactions $\pi N\to\pi N$ and $\pi N\to\pi\pi N$. As we have done an expansion in $M/m$ of the IR and EOMS expressions our results should in fact correspond to the one in the heavy baryon approach. Finally, $\bar{d}_{18}=-0.8\,$GeV-2 can be related to the Goldberger-Treiman discrepancy [38, 39]. For the LEC $C_{0}\equiv\sum_{i}C_{0}^{(i)}g_{0}^{i}$ we assume it to be of natural size, typically of the order of $1/\Lambda^{4}_{\chi}$ with $\Lambda_{\chi}=0.6\,$GeV an estimate of the breakdown scale of the chiral expansion (which is a more conservative estimate than given above and was used in Refs. [19, 37]). In Fig. 3 (left panel) we show the chiral expansion of $g_{A}$ at second, third and fourth order for set 1 of the LECs and similarly for set 2 in the right panel of Fig. 3. The upper fourth order curve corresponds to $C_{0}(m)=15\,$GeV-4, whereas the lower fourth order curve represents the case $C_{0}(m)=-15\,$GeV-4. While the third order correction is large, as the coefficient $\alpha_{3}$ in Eq. (1) is proportional to the large factor $2c_{4}-c_{3}$, we see that the fourth order corrections are rather small for pion masses below 300 MeV, even though some of the dimension-three LECs are rather large. Also, we find that $g_{0}=1.0$ for set 1 and $g_{0}=1.3$ for set 2, in order.
@@
Figure 3: $g_{A}$ as a function of $M$ for set 1 (left panel) and set 2 (right panel) of the LECs given in the text. The blue dash-dotted, the black dashed and the red solid lines represent the results up to $M^{2}$, $M^{3}$ and $M^{4}$, respectively. The upper $M^{4}$ curve corresponds to $C_{0}(m)=15\,$GeV-4, whereas the lower $M^{4}$ curve represents the case $C_{0}(m)=-15\,$GeV-4. The pink circle denotes the physical value of $g_{A}$.
@@
${\rm set}~{}1$ ::: $\Delta^{(2)}=1.5\%~{},\,~{}~{}~{}\quad\Delta^{(3)}=28.8\%~{},\quad\Delta^{(4)}=2.6\%~{},$
${\rm set}~{}2$ ::: $\Delta^{(2)}=-26.7\%~{},\quad\Delta^{(3)}=44.5\%~{},\quad\Delta^{(4)}=-17.4\%~{},$ (35)